# Identifying non-Hermitian critical points with quantum metric

### v1

@@ IV.2 Non-Hermitian mixed-field Ising model
where $h_{x}$ and $h_{z}$ denote the strengths of the real and imaginary fields that are along the transverse and longitudinal directions, respectively. The order parameter for the ground state $|\psi_{1}\rangle$ in this model is the magnetization $M_{z}=\frac{1}{N}\bra{\psi_{1}}\sum_{l=1}^{N}\sigma^{z}_{l}\ket{\psi_{1}}$, where |$\ket{\psi_{1}}$ can be numerically obtained by the exact diagonalization method for small system size $N$. The numerical results for $N=10$ under the PBC and the open boundary condition (OBC) are shown in Figs. 6(a-c) and Fig. 6(d), respectively. The phase diagram that contains the paramagnetic phase with $|M_{z}|\approx 0$ (and real energy) and the ferromagnetic phase with finite value of $|M_{z}|$ (and complex energy) in the $h_{z}$-$h_{x}$ parameter plane is shown in Fig. 6(a). Moreover, the phase transition between two phases is accompanied by a real-complex transition of the ground state energy. To detect the phase transition with the quantum metric, we numerically compute the logarithm quantify of quantum metric $\xi_{h_{z}}=\log_{10}g_{h_{z}h_{z}}$ as a function of $h_{z}$ and $h_{x}$ in Fig 6(b). One can see the boundary between the paramagnetic and ferromagnetic phases are well revealed by $\xi_{h_{z}}$. As shown in Fig. 6(c) and Fig. 6(d) for different boundary conditions with fixed $h_{x}=3$, the critical point $h_{z}\approx 0.9$ and $h_{z}\approx 0.96$ can both be clearly revealed by the quantum metric.
@@
Figure 7: (Color online) (a) $\xi_{V_{1}}$ as a function of $V_{1}$ for $H_{\text{GAA}}^{(1)}$ with different system sizes $L$ and fixed $g=h=0$ under periodic boundary condition. $\xi_{V_{1}}$ as a function of $V_{1}$ for $\tilde{H}$ [see Eq. (23)] with fixed (b) $g=0$; and (c) $g=0.5$, for different boundary condition coefficients $\zeta$. (d) The probability amplitudes of the ground-state wave functions as a function of $V_{1}$ under the PBC and the OBC, respectively. The system size in (b-d) is $L=610$. Other parameters are $h=0$, $V_{2}=0.5$, and $t=1$.
@@
Before concluding, we discuss the finite-size effect and boundary conditions on detecting the non-Hermitian critical points with the quantum metric. Without loss of generality, we focus on the non-Hermitian GAA model $H_{\text{GAA}}^{(1)}$ in Eq. (5). The finite-size scaling of the quantum metric of the ground state $\xi_{V_{1}}=\log_{10}g_{V_{1}V1}$ under the PBC for the lattice sizes $L=34,144,610,2584$ (chosen from the Fibonacci sequences) is shown in Fig 7(a). One can find that the critical point can be revealed by the peak of the quantum metric even for a small lattice of $L=34$, and the peak becomes higher as the lattice size is increased. To consider the boundary effect on the quantum metric, we add an boundary coupling term to the Hamiltonian $H_{\text{GAA}}^{(1)}$ as
$\begin{split}\tilde{H}=H_{\text{GAA}}^{(1)}+\xi(e^{-g}c^{\dagger}_{1}c_{L}+e^{g}c^{\dagger}_{L}c_{1}),\end{split}$ (23)
@@
where $\zeta$ is the coefficient for tuning different boundary conditions. For $\zeta=0$, it denotes the open boundary condition, while $\zeta=1$ ($\zeta\neq 0,1$) denotes the standard (modified) periodic boundary condition. As shown in Fig. 7(b), the results of $\xi_{V_{1}}$ as a function of $V_{1}$ under different boundary conditions ($\zeta=0,0.5,1$) coincide in the Hermitian limit. In the non-Hermitian case, the results of $\xi_{V_{1}}$ are different for periodic ($\zeta=0.5,1$) and open ($\zeta=0$) boundary conditions, while the transition point can revealed in both conditions, as shown in Fig 7(c). The difference comes from the localized bulk eigenstates induced by the non-Hermitian skin effect under the open boundary condition Yao and Wang (2018); Song et al. (2019), as shown in Fig 7(d). The non-Hermitian skin effect destroys the extended phase under the original periodic boundary condition and leads to the increase of the quantum metric with accidental degeneracies. Thus, our results are robust against to the finite-size effect and different boundary conditions when using the quantum metric to reveal critical points.
@@
In summary, we have investigated the quantum phase transitions in various non-Hermitian systems by using the information-geometric approach. It has been shown that the peaks of the quantum metric of the eigenstates exactly identify the localization transition points and mobility edges in the non-Hermitian GAA models. For non-Hermitian cluster and mixed-field Ising models, we have demonstrated that the phase boundaries in the non-Hermiticity parameter space determined by quantum metric of ground state perfectly coincide with those from the corresponding order parameters. These results indicate that the peak of the quantum metric serving as good signatures for detecting the non-Hermitian critical points in both single-particle and many-body systems. This strategy is robust against the finite-size effect and different boundary conditions. Notably, our present work only focuses on the right eigenstates of non-Hermitian Hamiltonians, and the quantum metric effect based on the biorthogonal eigenstates and the non-unitary dynamics are deserving to be further explored.

### v3

@@ IV.2 Non-Hermitian mixed-field Ising model
where $h_{x}$ and $h_{z}$ denote the strengths of the real and imaginary fields that are along the transverse and longitudinal directions, respectively. The order parameter for the ground state $|\psi_{1}\rangle$ in this model is the magnetization $M_{z}=\frac{1}{N}\bra{\psi_{1}}\sum_{l=1}^{N}\sigma^{z}_{l}\ket{\psi_{1}}$, where |$\ket{\psi_{1}}$ can be numerically obtained by the exact diagonalization method for small system size $N$. The numerical results for $N=10$ under the PBC and the open boundary condition (OBC) are shown in Figs. 6(a-c) and Fig. 6(d), respectively. The phase diagram that contains the paramagnetic phase with $|M_{z}|\approx 0$ (and real energy) and the ferromagnetic phase with finite value of $|M_{z}|$ (and complex energy) in the $h_{z}$-$h_{x}$ parameter plane is shown in Fig. 6(a). Moreover, the phase transition with gap closing between two phases is accompanied by a real-complex transition of the ground state energy. To detect the phase transition with the quantum metric, we numerically compute the logarithm quantify of quantum metric $\xi_{h_{z}}=\log_{10}g_{h_{z}h_{z}}$ as a function of $h_{z}$ and $h_{x}$ in Fig 6(b). One can see the boundary between the paramagnetic and ferromagnetic phases are well revealed by $\xi_{h_{z}}$. As shown in Fig. 6(c) and Fig. 6(d) for different boundary conditions with fixed $h_{x}=3$, the critical point $h_{z}\approx 0.9$ and $h_{z}\approx 0.96$ can both be clearly revealed by the quantum metric.
@@
Figure 7: (Color online) (a) $\xi_{V_{1}}$ as a function of $V_{1}$ for $H_{\text{GAA}}^{(1)}$ and (b) $\xi_{\Delta}$ as a function of $\Delta$ for $H_{\text{GAA}}^{(2)}$, with both $g=0.5$ and different system sizes $L$. (c) Finite-size scaling of $\xi_{V_{1c}}=\xi_{V_{1}}(V_{1}=V_{c})$ obtained from (a) with $V_{1c}=3.16$ and $\xi_{V_{1c}}=1.992\log_{10}L-2.983$. (d) Finite-size scaling of $\xi_{\Delta_{c}}=\xi_{\Delta}(\Delta=\Delta_{c})$ obtained from (b) with $\Delta_{c}=1.45$ and $\xi_{\Delta_{c}}=2.142\log_{10}L-2.708$. $\xi_{V_{1}}$ as a function of $V_{1}$ for $\tilde{H}$ [see Eq. (27)] with fixed (e) $g=0$ and (f) $g=0.5$, for different boundary condition coefficients $\zeta$. (g) $\eta$ as a function of $V_{1}$ for different boundary condition coefficients $\zeta$ with fixed $g=0.5$. (h) The density distribution of the ground-state with fixed $V_{1}=0$ and $g=0.5$ under the PBC and the OBC, respectively. The system size is $L=610$. Other parameters are $h=0$, $V_{2}=0.5$, $\alpha=-0.5$ and $t=1$.
@@
Before concluding, we discuss the finite-size effect and boundary conditions on detecting the non-Hermitian critical points with the quantum metric. Without loss of generality, we focus on the non-Hermitian GAA model $H_{\text{GAA}}^{(1)}$ in Eq. (5) and $H_{\text{GAA}}^{(2)}$ in Eq. (10). The finite-size scaling of the quantum metric of the ground state $\xi_{V_{1}}=\log_{10}g_{V_{1}V_{1}}$ for $H_{\text{GAA}}^{(1)}$ and $\xi_{\Delta}=\log_{10}g_{\Delta\Delta}$ for $H_{\text{GAA}}^{(2)}$ are shown in Figs. 7(a) and 7(b), respectively. The lattice sizes $L=34,144,610,2584$ are chosen from the Fibonacci sequences to satisfy the PBC. The critical points are determined as $V_{1c}\approx 3.16$ and $\Delta_{c}\approx 1.45$, respectively. We also plot the quantum metric at the critical points $\xi_{V_{1c}}=\xi_{V_{1}}(V_{1}=V_{c})$ and $\xi_{\Delta_{c}}=\xi_{\Delta}(\Delta=\Delta_{c})$ versus $L$ in Figs. 7(c) and 7(d), respectively. One can find that the quantum metric at the critical points exhibits the scaling law $g_{\mu\mu}\sim L^{\kappa}$ (i.e., $\xi_{\mu\mu}\sim\kappa\log_{10}L$ up to a constant), where the exponent $\kappa$ is related to the critical exponents Gu (2010); Lv et al. (2022a, b). The critical exponents for the non-Hermitian localization and quantum phase transitions may be extracted from the scaling analysis of the quantum metric and order parameters, which require further investigations.
To consider the boundary effect on the quantum metric, we add an boundary coupling term to the Hamiltonian $H_{\text{GAA}}^{(1)}$ as
$\begin{split}\tilde{H}=H_{\text{GAA}}^{(1)}+\zeta(e^{-g}c^{\dagger}_{1}c_{L}+e^{g}c^{\dagger}_{L}c_{1}),\end{split}$ (27)
@@
where $\zeta$ is the coefficient for tuning different boundary conditions. For $\zeta=0$, it denotes the open boundary condition, while $\zeta=1$ ($\zeta\neq 0,1$) denotes the standard (modified) periodic boundary condition. As shown in Fig. 7(e), the results of $\xi_{V_{1}}$ as a function of $V_{1}$ under different boundary conditions ($\zeta=0,0.5,1$) coincide in the Hermitian limit. In the non-Hermitian case, the results of $\xi_{V_{1}}$ are different for periodic ($\zeta=0.5,1$) and open ($\zeta=0$) boundary conditions, while the transition point can revealed in both conditions, as shown in Fig. 7(f). The difference comes from the localized bulk eigenstates induced by the non-Hermitian skin effect under the open boundary condition Yao and Wang (2018); Song et al. (2019), as shown in Figs. 7(g) and (h). The non-Hermitian skin effect destroys the extended phase under the original periodic boundary condition and leads to the increase of the quantum metric with accidental degeneracies. Thus, our results are robust against to the finite-size effect and different boundary conditions when using the quantum metric to reveal critical points.
@@
In summary, we have investigated the quantum phase transitions in various non-Hermitian systems by using the information-geometric approach. It has been shown that the peaks of the quantum metric of the eigenstates exactly identify the localization transition points and mobility edges in the non-Hermitian GAA models. For non-Hermitian cluster and mixed-field Ising models, we have demonstrated that the phase boundaries in the non-Hermiticity parameter space determined by quantum metric of ground state perfectly coincide with those from the corresponding order parameters. These results indicate that the peak of the quantum metric serving as good signatures for detecting the non-Hermitian critical points with gap closings in both single-particle quasi-periodic and many-body spin systems. This strategy is robust against the finite-size effect and different boundary conditions. Notably, our present work only focuses on the right eigenstates of non-Hermitian Hamiltonians, and the quantum metric effect based on the biorthogonal eigenstates and the non-unitary dynamics are deserving to be further explored.
-Note added. After finishing our manuscript, we noticed two relevant preprints Orlov et al. ; Zeng et al. . In Ref. Orlov et al. , a generalization of the QGT based on the generator of adiabatic transformations was developed to study phase transitions in non-Hermitian quantum systems. In Ref. Zeng et al. , the self-normal and biorthogonal fidelity susceptibilities were used to study the critical behaviors in the nonreciprocal Aubry-André-Harper model.